OpenStudy (anonymous):

When x is divided by 4, the remainder is 3. What is the remainder when 2x is divided by 4?

OpenStudy (anonymous):

Is it 2 ?

OpenStudy (anonymous):

Try expressing the number x in terms of a quotient Q and the remainders.

OpenStudy (lgbasallote):

you do know 7 divided by 4 has remainder 3 too right @ParthKohli

OpenStudy (anonymous):

Like x = 4Q + 3. Now 2x = 8Q + 6 2x = 8Q + 4 +2 = 4(2Q + 1 ) + 2

OpenStudy (anonymous):

Where Q is an integer. :)

OpenStudy (anonymous):

solution is above, but a trick for these problems is to try it with numbers and see what you get

OpenStudy (anonymous):

try it with 7 14 divided by 4 leaves a remainder of 2 or try it with 11 22 divided by 4 leaves a remainder of 2 should be clear right?

OpenStudy (anonymous):

Yes :) Thanks haha :)